manu is 4 years older than his sister anu. if the sum of their ages is 20 years, find their ages.
Answer by homeworkhedgehog(9)

For this, you can create two equations and solve them using substitution or elimination.
"Manu is 4 years older than his sister Anu" translates to
M = 4 + A
"The sum of their ages is 20 years" translates to
M + A = 20
Since we know that M = 4 + A, we can substitute that into "M" in the other equation:
(4 + A) + A = 20
Now we combine our "like" terms (the parentheses are negligible):
4 + 2A = 20
And solve for "A" using algebra.
Subtract 4 from both sides:
2A = 16
Divide both sides by 2:
A = 8
Now we know how old Anu is. She is 8 years old. We use this to find Manu's age. "Manu is 4 years older than Anu" tells us that Manu is 12. This checks out when we add 8 and 12 to see if their combined age is really 20 like the problem says.